# Impurities in a trapped 1D Bose gas of arbitrary interaction strength: localization-delocalization transition and absence of self-localization

### v1

@@
We discuss impurities in a trapped one-dimensional Bose gas with arbitrary boson-boson and boson-impurity interactions. To fully account for quantum effects, in particular in the regime of strong boson-boson interactions, we employ numerical simulations based on the density-matrix renormalization group (DMRG) and analytic approximations exploiting the mapping to weakly interacting fermions. Mobile impurities in a box potential undergo a phase transition between a delocalized state and a solution localized at one of the potential edges upon increasing the impurity-boson interaction. While a mean-field ansatz based on coupled Gross-Pitaevski – Schrödinger equations gives reasonable predictions of this transition, it also predicts the existence of a self-localized polaron solution, which we show to be an artifact of the underlying decoupling approximation. This demonstrates that impurity-boson correlations are important even in the limit of weak boson-boson interactions, where mean-field approaches are expected to work well. Furthermore we calculate the energy of a single polaron formed by a heavy impurity for arbitrary interaction strengths and give analytical approximations for large but finite boson-boson couplings. Finally we numerically determine the polaron-polaron interaction potential in Born-Oppenheimer approximation, which in the Tonks gas limit is oscillatory due to Friedel oscillations in the Bose gas.
@@ I Introduction
Quasi particles formed by quantum impurities immersed in a many-body environment play a key role for the understanding of transport phenomena [1]. Furthermore the interaction of quasi particles mediated by the host medium forms the basis of many important many-body phenomena in condensed-matter physics [2, 3, 4, 5]. In recent years the possibility to experimentally investigate impurities in quantum fluids, such as Bose-Einstein condensates (BEC) of atoms has renewed the interest in these quasi-particles, called Bose polarons. An important difference between Bose polarons in ultra-cold quantum gases and the Landau-Pekar polaron [6, 7] introduced to model electrons interacting with the lattice vibrations of a solid is the large compressibility of the BEC. As a consequence the standard model for polarons in solids, the Fröhlich model [8], which describes the polaron as interaction of the impurity with phonon excitations, is only suitable for weak boson-boson and weak impurity-boson interactions [9]. In this limit the condensate can be considered undepleted and the role of lattice vibrations is taken over by the Bogoliubov phonons.
@@
For stronger interactions the back-action of the impurity to the condensate needs to be taken into account, while keeping correlations between impurity and bosons. For translational invariant systems this can be done by means of a Lee-Low-Pines (LLP) transformation [10], which decouples the impurity motion from the many-body problem of interacting bosons. For weak boson-boson interactions the latter can then be treated rather accurately in a mean-field description for arbitrary strength of the impurity-boson coupling [11, 12, 13, 14, 15].
@@
In an inhomogeneous Bose gas and for repulsive impurity-boson interactions, the polaron can localize in density minima of the Bose gas and, similarly to phase separation in multi-component condensates [16], a localization-delocalization transition can occur. Due to the absence of translational invariance the LLP decoupling does not work here, however. Instead another mean-field ansatz is often used, which in addition neglects correlations between impurity and bosons and results in a coupled Gross-Pitaevski and Schrödinger equation for the condensate and the impurity respectively [17, 18, 19, 20, 21, 22]. This decoupling mean-field approach (DMF) however also predicts a localization transition to occur in translationally invariant systems. In analogy to impurities in liquid 4He [23, 24], a single atom can become self-trapped in a localized distortion of the condensate created by the impurity [17, 18, 19, 20, 21, 22] thereby spontaneously breaking translational invariance. In 2D and 3D a critical interaction strength is needed for such a self-trapped polaron to exist, but in 1D arbitrarily small impurity-boson interactions are sufficient [21, 22].
@@
In the present paper we analyze the role of quantum fluctuations to localization-delocalization transitions of impurities in BECs. To this end we consider a mobile impurity of finite mass $M$ in a one-dimensional condensate and employ numerical simulations based on the density-matrix renormalization group (DMRG) [25, 26]. Furthermore we consider a box potential, see Fig. 1, to explicitly break translational invariance. For such a system decoupling mean-field theory predicts a reentry transition from a self-localized polaron at the trap center to a polaron localized at the potential edge upon changing the impurity-boson coupling strength. We numerically determine the ground-state phase diagram taking quantum fluctuations into account and observe instead only a phase transition between a fully delocalized phase for small values of the impurity-boson coupling $g_{\text{IB}}$ and a localized phase near one of the edges. While the latter is in qualitative agreement with DMF predictions, our simulations show that there is no self-trapped solution in the full quantum problem.
We proceed by considering a heavy impurity $M\gg m$ and calculate the polaron energy as function of impurity-boson coupling for different boson-boson interaction strength ranging from the Bogoliubov regime to the Tonks-gas regime of (nearly) impenetrable bosons. Making use of the mapping between strongly interacting bosons and weakly interacting fermions, we derive analytical approximations of the polaron energy which agree very well with the numerical DMRG data. Finally we calculate the effective interaction potential between two heavy impurities in Born-Oppenheimer approximation for short distances. In a weakly interacting Bose gas the potential is monotonous and has a linear slope for small distances [13]. In the Tonks gas limit of impenetrable bosons the potential is modified by oscillations due to Friedel-like density oscillations of the hard-core bosons.
The paper is organized as follows: In Sec. II we introduce the model and briefly discuss the numerical methods used. Then we discuss a finite-mass impurity in a one-dimensional Bose gas in a trap potential in Sec. III. Finally we discuss the properties of a single polaron formed by a heavy impurity in a strongly interacting Bose gas in Sec. IV and the interaction between two heavy polarons in Born-Oppenheimer approximation in Sec. V.
II Model and numerical method
II.1 Model
@@
where $n_{0}$ is the mean particle density of bosons in the system. The Tonks parameter gives a relation between the interaction ($\propto gn_{0}^{2}$) and kinetic energies ($\propto n_{0}^{3}/m$) of the bosons. A large value of $\gamma$ corresponds to strong boson-boson interactions. For $\gamma\ll 1$ and below a critical temperature a quasi condensate forms. (True condensation is not possible due to the Mermin-Wagner-Hohenberg theorem [27, 28].) For $\gamma\rightarrow\infty$ the bosons become impenetrable hard-core bosons and can be mapped to free fermions [29].
II.2 DMRG simulation of quantum impurity in a 1D Bose gas
@@
A powerful method to describe the ground-state of quantum many-body systems in 1D is the density matrix renormalization group (DMRG) [25]. The DMRG method was originally developed for lattice systems. It can also be applied to continuous models, however, e.g. using a proper discretization, which we apply here. To this end let us first look at the Hamiltonian where both bosons and the impurity are described in second quantization:

### v2

@@
We discuss impurities in a one-dimensional Bose gas with arbitrary boson-boson and boson-impurity interactions. To fully account for quantum effects, we employ numerical simulations based on the density-matrix renormalization group (DMRG) and - in the regime of strong boson-boson interactions - the mapping to weakly interacting fermions. A mean-field description of the Bose polaron based on coupled Gross-Pitaevski – Schrödinger equations predicts the existence of a self-localized polaron. We here show that such a solution does not exist and is an artifact of the underlying decoupling approximation. To this end we consider a mobile impurity in a box potential. Our work demonstrates that correlations between the impurity position and the bosons are important even in the limit where mean-field approaches are expected to work well. Furthermore we derive analytical approximations for the energy of a single polaron formed by a heavy impurity for arbitrary interaction strengths and large but finite boson-boson couplings which accurately reproduce DMRG results. This demonstrates that the polaron problem of a heavy impurity in a 1D Bose gas can be accurately approximated by a proper mean-field description plus a linearized treatment of quantum fluctuations for arbitrary boson-boson and impurity-boson couplings. Finally we determine the polaron-polaron interaction potential $V(r)$ in Born-Oppenheimer approximation for small and intermediate distances $r$, which in the Tonks gas limit is oscillatory due to Friedel oscillations in the Bose gas.
@@ I Introduction
Quasi particles formed by quantum impurities immersed in a many-body environment play a key role for the understanding of transport phenomena [1]. Furthermore the interaction of quasi particles mediated by the host medium forms the basis of many important many-body phenomena in condensed-matter physics [2, 3, 4, 5]. In recent years the possibility to experimentally investigate impurities in quantum fluids, such as Bose-Einstein condensates (BEC) of atoms has renewed the interest in these quasi-particles, called Bose polarons. An important difference between Bose polarons in ultra-cold quantum gases and the Landau-Pekar polaron [6, 7] introduced to model electrons interacting with the lattice vibrations of a solid is the large compressibility of the BEC. As a consequence a common theoretical model similar to the Fröhlich model in solids [8], which describes the polaron as interaction of the impurity with phonon excitations, is only suitable for weak boson-boson and weak impurity-boson interactions [9]. In this limit the condensate can be considered undepleted and the role of lattice vibrations is taken over by the Bogoliubov phonons.
@@
For stronger interactions with the impurity the back-action to the condensate needs to be taken into account, while keeping correlations between impurity position and bosons. For translational invariant systems this can be done by means of a Lee-Low-Pines (LLP) transformation [10], which decouples the impurity motion from the many-body problem of interacting bosons. For weak boson-boson interactions $g$ the latter can then be treated rather accurately in a mean-field description for arbitrary strength of the impurity-boson couplings $g_{\textrm{IB}}$ [11, 12, 13, 14, 15].
@@
In an inhomogeneous Bose gas and for repulsive impurity-boson interactions, the polaron can localize in density minima of the Bose gas and, similarly to phase separation in multi-component condensates [16], a localization-delocalization transition can occur. Due to the absence of translational invariance the LLP decoupling does not work here, however. Instead another mean-field ansatz is often used, which in addition neglects however correlations between impurity position and bosons and results in a coupled Gross-Pitaevski and Schrödinger equation for the condensate and the impurity respectively [17, 18, 19, 20, 21, 22]. This decoupling mean-field approach (DMF) also predicts a localization transition to occur in translationally invariant systems. In analogy to impurities in liquid 4He [23, 24], a single atom can become self-trapped in a distortion of the condensate created by the impurity [17, 18, 19, 20, 21, 22] thereby spontaneously breaking translational invariance. In 2D and 3D a critical interaction strength is needed for such a self-trapped polaron to exist, but in 1D within the DMF approach an arbitrarily small $g_{\text{IB}}$ is sufficient [21, 22]. Recently it was shown in Ref.[25] that correlations between the center-of-mass motion of the impurity and the bosons suppress this self-localization transition and solutions with finite localization length only survive for large $g_{\text{IB}}$.
@@
In the present paper we analyze the role of quantum fluctuations to the self-trapping. To this end we consider a mobile impurity of finite mass $M$ in a one-dimensional condensate and employ numerical simulations based on the density-matrix renormalization group (DMRG) [26, 27]. Furthermore we consider a box potential, see Fig. 1, to explicitly break translational invariance. For such a system a DMF approach predicts a reentry transition from a self-localized polaron at the trap center to a delocalized polaron and eventually to one localized at the potential edge upon changing $g_{\text{IB}}$. We numerically determine the ground-state phase diagram taking quantum fluctuations into account and observe instead only a phase transition between a fully delocalized phase for small values of the impurity-boson coupling $g_{\text{IB}}$ and a localized phase near one of the edges. Our simulations show that there is no self-trapped solution in the full quantum problem even for very large impurity-boson couplings for which a finite localization length was obtained in Ref.[25].
In [12, 13, 14] we have shown that for a heavy impurity with mass ratios $M/m>{\cal O}(1)$ in a translational invariant, weakly interacting 1D Bose gas, a mean-field approach in a LLP frame gives very accurate predictions for most properties of the polaron, which can further be improved by adding quantum fluctuations perturbatively. We here show that this is also true for a stronly-interacting 1D Bose gas. To this end we calculate the polaron energy as function of $g_{\text{IB}}$ for arbitray boson-boson interaction strength ranging from the Bogoliubov regime to the Tonks-gas regime of (nearly) impenetrable bosons. Making use of the mapping between strongly interacting bosons and weakly interacting fermions, we derive an analytical approximation of the polaron energy in the latter regime which agrees very well with the numerical DMRG data.
Finally we calculate the effective interaction potential between two heavy impurities in Born-Oppenheimer approximation for short distances. In a weakly interacting Bose gas the potential is monotonous and has a linear slope for small distances [13]. In the limit of (nearly) impenetrable bosons the potential is modified by Friedel-like oscillations, whose long-range behavior has been derived in [28, 29] using a low-energy approximation. We here extend this approach to also capture the short-distance behavior, relevant for bi-polaron bound states.
The paper is organized as follows: In Sec. II we introduce the model. We discuss the problem of self-localization of a finite-mass impurity in a one-dimensional Bose gas in a box potential in Sec. III. Then we discuss a mean-field approach to single polaron formed by a heavy impurity in a strongly interacting Bose gas in Sec. IV. Finally we derive the short-range interaction potential between two heavy polarons in Born-Oppenheimer approximation in Sec. V.
II Model
@@
where $n_{0}$ is the mean particle density of bosons in the system. The Tonks parameter gives a relation between the interaction ($\propto gn_{0}^{2}$) and the kinetic energy ($\propto n_{0}^{3}/m$) of the bosons. A large value of $\gamma$ corresponds to strong boson-boson interactions. For $\gamma\ll 1$ and below a critical temperature a quasi condensate forms. (True condensation is not possible due to the Mermin-Wagner-Hohenberg theorem [30, 31].) For $\gamma\rightarrow\infty$ the bosons become impenetrable hard-core bosons and can be mapped to free fermions [32].
@@
A powerful method to describe the ground-state of quantum many-body systems in 1D is the density matrix renormalization group (DMRG) [26]. The DMRG method was originally developed for lattice systems. It can also be applied to continuous models, however, e.g. using a proper discretization, which we apply here. The details of this together with benchmarks are given in the Appendix A.
$\quad+g|\phi(x,t)|^{2}+V_{B}(x)\biggr{)}\phi(x,t),$ (3)
We will show here that in contrast to a mean-field theory in the LLP frame, this decoupling approach leads to artifacts.
III Localization-Delocalization transition in a box potential
In [33, 34] it has been argued based on a DMF ansatz, eq. (3), that in a translational invariant, homogeneous system the impurity wavefunction should self-trap in a co-localized distortion of the BEC for small $g_{\text{IB}}$. The reality of such self-trapped polarons has been subject of discussions. Recently it was shown in Ref. [25] that taking into account leading-order correlations, self trapping only occurs above a certain minimum value of $g_{\text{IB}}$.
We now show that in a full quantum model self trapping is absent altogether. To this end we numerically investigate a box potential for bosons and impurity $V_{B}(x)=V_{I}(x)$. Here we expect a transition between a delocalized phase, where the impurity is spread out over the whole system, and a localized one, where the impurity is localized at one of the two edges of the system. This is because the energy of an impurity at the edge increases as $E_{\textrm{loc}}\sim\sqrt{g_{\text{IB}}}$, while the polaron energy in the bulk scales as $E_{p}\sim g_{\text{IB}}$. Thus for $g_{\text{IB}}\to 0$ no edge state exists which only emerges above a critical value of $g_{\text{IB}}$. In addition to this a decoupling mean-field theory predicts another phase transition to a self-localized polaron. Due to the broken translational invariance in the box potential considered here, such a self-trapping would manifest itself in a probability density of the impurity centered in the middle of the trap with a width independent on the trap size $L$.
III.1 Localization-delocalization transition in a box potential
We performed DMRG simulations of the ground state of a mobile impurity in a 1D Bose gas in the box potential for different mass ratios $M/m$ and impurity-boson interaction strengths $g_{\text{IB}}$. Results for $\gamma=0.4$, and $\gamma=\infty$ are shown in Fig. 2 and Fig. 3, respectively. One recognizes in both cases a sharp transition between a phase where the impurity is delocalized over the trap to a phase where it is localized at one of the two edges. (Note that the true ground state is a superposition of localized states at both edges. Since the energy difference between the symmetric and antisymmetric superpositions vanishes exponentially with increasing system size, the DMRG algorithm converges to a solution on one edge only.) The transition point depends on the boson-boson interaction strength. In the delocalized phase the probability density of the impurity is spread out over the whole system and apart from some corrections at the edges it corresponds to the ground state of the box potential $\sim\cos^{2}(x\pi/L)$, with $L$ being the length of the box. In the edge-localized phase, on the other hand, the width of the probability distribution becomes independent on system size for large $L$.
Abbildung 2: a) Localization-delocalization transition of impurity in a box potential in a weakly interacting Bose gas with Tonks parameter $\gamma=0.4$ and $N=40$ particles. Color code describes average position $\langle\hat{X}\rangle$ of impurity. Solid line is the prediction from eq. (6) with $l_{0}=\xi$. The dashed line gives the prediction with $l_{0}$ from a variational ansatz. b) and c) density distribution corresponding to the two points in Figure a).
$E_{p}\approx g_{\text{IB}}n_{0}$ (4)
and thus scales linearly in $g_{\text{IB}}$.
For a weakly interacting Bose gas (small $\gamma$) the characteristic length scale $l_{0}=\eta\xi$ of the harmonic confinement is the healing length $\xi$ of the Bose condensate. The factor $\eta\simeq{\cal O}(1)$ accounts for the backaction of the trapped impurity on the density of the condensate near the potential edge, which is relevant in the small-$\gamma$ regime, due to the large compressibility of bose gas. The impurity thus experiences an effective potential for small $x$:
$\phi(x)=\sqrt{n_{0}}\tanh\left(\alpha x\right).$ (7)
describing the coherent amplitude of the condensate. $|\phi_{I}\rangle$ is then the ground state of the impurity in the resulting effective Pöschl-Teller potential, which can be calculated analytically [35]. This approach has already been used to determine bound states of the impurity in a box potential using the ansatz $\alpha=1/(\sqrt{2}\xi)$ [36]. Here we instead determine the characteristic length scale $\alpha$ by minimizing the total energy. This gives the implicit equation for $\eta=1/(\sqrt{2}\xi\alpha)$ as function of the mass ratio $M/m$, the Tonks parameter $\gamma$ and $\gamma_{\textrm{IB}}=g_{\text{IB}}m/n_{0}$
Abbildung 3: a) Localization-delocalization transition of an impurity in a box potential in a Tonks gas ($\gamma=\infty$, $N=40$). Solid line is prediction from eq. (10). b) and c) density distribution corresponding to the two points in Figure a).
$n(x)\approx\frac{2k_{F}}{\pi}\left(1-\frac{\sin(2k_{F}x)}{2k_{F}x}\right),$ (9)
with $x$ being the distance from the edge and $k_{F}=\pi n_{0}/2$ is the Fermi momentum. The characteristic length scale $l_{0}$ is here $(\sqrt{6}/\pi)n_{0}^{-1}$ which gives the following estimate for the critical $g_{\text{IB}}$ of the localization-delocalization transition:
Abbildung 4: Critical $g_{\text{IB}}$ of the localization-delocalization transition for different Tonks parameters $\gamma$, a constant mass ratio $M/m=2$ and $N=40$ bosons calculated from DMRG simulations (crosses). Also shown is the critical $g_{\text{IB}}$ derived from DMRG simulations of a Tonks gas $\gamma\rightarrow\infty$ (solid blue line). These values are compared to the predictions for a weak interacting Bose gas made by the variational ansatz (orange dashed line) and predictions for a Tonks gas eq. (10) (green dashed line).
In Fig. 4 we compare $g_{\text{IB}}|_{\textrm{crit}}$, obtained from DMRG simulations with the analytic predictions for small, eq. (6), and large Tonks parameters $\gamma$, eq. (10). Since the box potential is finite, we determine the transition point of the localization-delocalization transition by calculating at the uncertainty of the impurity position as function of $g_{\text{IB}}$. At the critical value $g_{\text{IB}}|_{\textrm{crit}}$ the position uncertainty is maximal. Analytical and numerical values align well up to a $\gamma\approx 2.5$, above this value the variational ansatz fails to produce a real value of $\eta$. For $\gamma>10$ the numerical values overshoot the analytic predictions from eq. (10). This is because we approximated the potential produced by the Tonks gas as a harmonic potential.
III.2 Absence of self-localization in a weakly interacting Bose gas
Abbildung 5: Comparison between average impurity position $\expectationvalue{\hat{X}}$ from DMRG and DMF calculations for $\gamma=0.4$ and $N=40$. a) Colorplot of $\expectationvalue{\hat{X}}$ as function of ${M}/{m}$ and $g_{\text{IB}}$ obtained by DMRG. Red lines show cuts which are plotted in c). b) The same obtained from DMF. c) Cuts in a) for two mass ratios, solid and dashed line correspond to each other. One clearly recognizes the absence of a self-trapped solution in the DMRG simulations.
Abbildung 6: Comparison between spatial variance of impurity $\langle\Delta\hat{X}\rangle$ from DMRG and DMF for $\gamma=0.4$ and $N=40$. a) Colorplot of variance from DMRG simulations as function of ${M}/{m}$ and $g_{\text{IB}}$. Red lines show cuts which are plotted in c). b) The same obtained from DMF. c) Cuts in a) for two mass ratios, solid and dashed line correspond to each other. The grey dashed-dotted line shows the spatial width of the impurity expected for $g_{\text{IB}}=0$. The black dotted line shows the self-localization length $\lambda$ predicted in DMF for $\frac{M}{m}=1.11$.
Abbildung 7: Left: Density distribution of Bose gas and impurity for different $g_{\text{IB}}$ (given in the lower right) from DMRG simulations. Right: The same from mean-field calculations compared to the analytic prediction for a self-trapped polaron, eq. (11). Parameters of the bose gas are $\gamma=0.4$, $M/m=1$ and $N=40$.
[…]
In summary the self-localisation is an artifact of the decoupling approximation and is not present in the full quantum approach. This is because boson-impurity correlations, which are neglected in DMF, suppress the self-localization [25]. Different from [25] self localization does not occur in the full quantum simulations even for large impurity-boson couplings.
IV Heavy polaron in a bose gas with arbitrary boson-boson interactions
In [12, 13] we have shown that key properties of heavy polarons in 1D Bose gases can be rather accurately be described by a mean-field theory in a LLP frame that takes into account the backaction of the impurity to the condensate. Quantum fluctuations have to be included only in a linearized approximation. These studies were however limited to weak interactions of the bosons, characterized by Tonks parameter $\gamma<1$. In the following we show that the opposite case of large Tonks parameters $\gamma\gg 1$ can be accurately treated as well by a perturbation expansion of a complementary mean-field description in the Tonks gas limit.
A characteristic quantity of the polaron is the energy $E_{\mathrm{P}}$ needed to immerse an impurity into the Bose gas $E_{\mathrm{P}}=E(g_{\text{IB}}\neq 0)-E(g_{\text{IB}}=0)$. Here $E(g_{\text{IB}}=0)$ is the ground state energy of the system without Bose-impurity-interaction. In the following we will compare analytic and semi-analytic predictions from linearized fluctuation expansions around mean field approaches in the weak- ($\gamma<1$) and strong-interaction cases ($\gamma\gg 1$) with exact numerical predictions. To this end we will employ DMRG simulations, which agree with previous exact results obtained by Quantum Monte Carlo simulations in Ref.[37].
IV.1 Polaron energy in a weakly interacting Bose gas
IV.2 Polaron energy in a strongly-interacting Bose gas
In the following we will discuss the case of a strongly interacting Bose gas, $\gamma\gg 1$. To this end we consider an impurity whose mass is much larger then that of the bosons ($\frac{M}{m}\gg 1$) such that the kinetic energy of the impurity can be neglected. The resulting effective Hamiltonian then reduces to a interacting boson problem in an external $\delta$ potential
Figure 8 shows the polaron energy in a strongly ($\gamma=20)$ interacting, trapped 1D Bose gas. The boson density shows Friedel like oscillations, see also [37], and the characteristic size of the condensate depletion around the fixed position of the impurity is given by the wavelength of these oscillations.
Abbildung 8: Polaron in strongly interacting 1D Bose gas with $\gamma=20$ and $N=80$ particles in a finite box and $M=\infty$: a) polaron energy as function of impurity-boson interaction, b), c) wavefunction for two different values of $g_{\text{IB}}$ (written in the lower right) in a box potential and impurity fixed at $x=0$.
IV.2.1 Tonks limit $\gamma\to\infty$
In the limit $\gamma\to\infty$ the 1D interacting Bose gas maps to free fermions. Here we can calculate the single particle solutions of free fermions with a infinitely heavy impurity at $x=0$ for open boundary conditions (OBCs). This system is described by the Hamiltonian
$\hat{H}_{F}^{0}=\int_{-L/2}^{L/2}\!\!\!\!\text{d}x\,\,\hat{\psi}^{\dagger}(x)\left(-\frac{\partial^{2}_{x}}{2m}+g_{\text{IB}}\delta\left(x\right)\right)\hat{\psi}(x),$ (14)
with $\{\hat{\Psi}(x),\hat{\Psi}^{\dagger}(y)\}=\delta(x-y)$. This means we need to solve the problem of non-interacting fermions trapped in a infinitely deep box potential (OBC’s) with a delta potential in the middle.
$E=\sum_{l=1}^{N}\frac{k_{l}^{2}}{2m},$ (16)
where the single particle solutions are given by
$\varphi_{l}(x)=a_{l}\mathrm{sgn}(x)^{l}\sin\left(k_{l}\left(|x|-\frac{L}{2}\right)\right).$ (17)
$l$ is an integer and the wave numbers $k_{l}$ have to be numerically calculated from the boundary condition at the delta potential
This then gives for the polaron energy, see [37]
$E_{p}$ $=$ $\frac{n_{0}^{2}\pi^{2}}{2m}\varepsilon(g_{\text{IB}}),$
$\varepsilon(g_{\text{IB}})$ $=$ $\frac{1}{\pi}\left[\left(1+\frac{\eta^{2}}{\pi^{2}}\right)\arctan\frac{\eta}{\pi}+\frac{\eta}{\pi}-\frac{\eta^{2}}{2\pi}\right],$ (19)
where $\eta=g_{\text{IB}}m/n_{0}>0$.
IV.2.2 Analytic approximation of the polaron energy for $1\ll\gamma<\infty$.
$\hat{H}_{F}=\hat{H}_{F}^{0}+\hat{H}_{1}$ (20)
with
$\hat{H}_{1}=-\frac{g_{\mathrm{F}}}{2}\int\!\!\!\int\text{d}x\,\text{d}y\,\,\hat{\psi}^{\dagger}(x)\hat{\psi}^{\dagger}(y)\overset{\leftarrow}{\frac{\partial}{\partial z}}\delta(z)\overset{\rightarrow}{\frac{\partial}{\partial z}}\hat{\psi}(y)\hat{\psi}(x),$ (21)
$g_{\mathrm{F}}=-\frac{4}{g}.$ (22)
The boson-fermion mapping provides an elegant way to find approximate analytic expressions for the polaron energy in a strongly interacting Bose gas with $\gamma\gg 1$ perturbatively in $|g_{\mathrm{F}}|\sim 1/\gamma$.
[…]
$\Delta E_{1}=\bra{\psi}\hat{H}_{1}\ket{\psi}$ (23)
Fig. 9 shows a comparison of the polaron energies as function of $g_{\text{IB}}$ from first order perturbation in $1/\gamma$, as well as from eq. (12) for small values of $\gamma$, with numerical DMRG results. In both cases the polaron energy saturates for increasing $g_{\text{IB}}$, which happens when the density of the Bose gas at the position of the impurity approaches zero. One recognizes rather good agreement between analytic approximations and exact simulations both for small and large values of $\gamma$ even down to $\gamma=20$ and for all values of $g_{\text{IB}}$. Note that the analytic approximation for $\gamma=\infty$ is that from Ref.[37].
Although we have considered here only the case of an impurity with an infinite mass and large values of $\gamma\geq 10$ and $\gamma\leq 0.1$ one recognizes that key properties of polarons formed by heavy impurities in a 1D Bose gas with arbitrary boson-boson and arbitrary boson-impurity interactions can be obtained rather accurately from a proper mean-field approach and potentially including lowest-order quantum corrections in Bogoliubov approximation.
For a weakly interacting Bose gas $(\gamma\ll 1)$ a mean-field ansatz in the LLP frame can be used to obtain semi-analytic expressions of the interaction potential at short distances, which agree very well with quantum Monte-Carlo simulations [13]. The mean field approach does not describe the Casimir-like contributions arizing from the exchange of virtual phonons [28, 29, 48, 49, 50, 51], which give however only important corrections in the tails of the interaction potential.
In this section we investigate the interaction between Bose polarons in a strongly-interacting 1D quasi condensate. In the Tonks limit $\gamma\to\infty$, where the interacting bosons can be mapped to free fermions, an analytic approximation to the interaction potential has been obtained in Ref.[28, 29] using a low-energy Luttinger-liquid approximation
$V(r)\approx\frac{v_{F}}{2\pi r}\textrm{Re}\,\textrm{Li}_{2}\left(-\frac{g_{\text{IB}}^{2}e^{2\pi in_{0}r}}{\bigl{(}v_{F}+ig_{\text{IB}}\bigr{)}^{2}}\right),\quad r\gg n^{-1}$ (25)
with $v_{F}=\pi n_{0}/m$ being the Fermi velocity, and $\textrm{Li}_{2}$ the dilogarithmic function. One notices an oscillatory behavior with the frequency being that of Friedel oscillations. The low-energy approximation gives an accurate description of the large distance behavior of the interaction potential including the Casimir contributions, but fails at short distances, which are however important for the formation of bi-polaron bound states. Thus we here determine the interaction potential both numerically by DMRG simulations and analytically in the Tonks-gas limit for all distances $r$. Since DMRG is not well suited for periodic boundary conditions, we again assume a confinement of the Bose gas to a box potential. As long as the distance of the impurities from the edges of the box is much larger than the healing length or the Fermi wavelength, boundary effects can be neglected. For the same reason our simulations, although in principle suitable, do not allow to accurately extract the far tails of the interaction potential, dominated by virtual phonon exchange. As discussed in Appendix B, the DMRG simulations reproduce the semi-analytical results from Ref.[13] obtained in mean-field theory for small values of $\gamma$.
$V(r)=E_{lr}(r)-E_{l}(r)-E_{r}(r)+E_{0}.$ (26)
Abbildung 10: Left: Interaction potential of heavy impurities in a bose gas with $\gamma=4$ and $N=100$ particles and bose-impurity interaction strength $g_{\text{IB}}m/n_{0}=1.08$. Right: Density of Bose gas at the two separations indicated in left picture as crosses in the corresponding color.
Abbildung 11: Left: Interaction potential of heavy impurities in the Tonks-gas limit $\gamma=\infty$, $N=80$ and $g_{\textrm{IB}}m/n_{0}=1.58$ The dashed line shows the Luttinger-liquid approximation from Ref.[28], eq. (25). Insert illustrates origin of oscillatory potential. Also shown is the analytic prediction (black line) of the interaction potential obtained from mapping to free fermions in PBC. Right: Density of Bose gas at the two separations indicated in left picture as crosses in the corresponding color.
For large values of $\gamma$ on the other hand oscillations appear in the polaron potential, as can be seen in Fig. 11. The frequency of these oscillations is that of the Friedel oscillations caused by the polarons themselves. As shown in Fig. 11b.), once a single impurity is introduced into the system it causes density oscillation and the second impurity has to displace a smaller or a larger amount of bosons and therefore needs less or more energy to form a polaron. The resulting energy modulations reflect themselves in the polaron potential. Also shown is a comparison to the analytic low-energy approximation, eq. (25), which diverges as $r\to 0$.
$V(r)=E_{\mathrm{lr}}-2E_{1}+E_{0}.$ (28)
As can be seen in Fig. 11 the bipolaron potential as well as the density distribution of the Tonks gas obtained in this way compare quite well with the DMRG simulations. These results are also in agreement with earlier findings in [52].
VI Summary
In the present paper we studied the ground state of a single and two impurities in a one-dimensional Bose gas for arbitrary impurity-boson and boson-boson interactions, addressing: (i) the existence of self-localization of polarons, (ii) the accuracy of mean-field descriptions of polarons which take impurity-boson correlations into account, and (iii) the Born-Oppenheimer interaction potential between two polarons beyond a low-energy approximation. To fully account for quantum effects within the Bose gas, which are particularly important in the limit of large Tonks parameters $\gamma$, we performed numerical simulations of a discretized effective lattice model using the density-matrix renormalization group.
In a commonly used mean-field approach to the Bose polaron, condensate and impurity are described by a factorized impurity-boson c-number wavefunction, leading to a coupled Gross-Pitaevski – Schrödinger equation. Such an approach predicts the existence of a self-trapped polaron for arbitrarily small impurity-boson couplings in a homogeneous 1D gas, where the center of mass of the impurity is localized in a distortion of the condensate created by the impurity. Such a decoupling mean-field (DMF) theory neglects spatial correlations between impurity and bosons. Recent findings have shown that including these correlations in leading order prevents self-trapping for small impurity-boson couplings [25]. We here showed by comparison with exact DMRG results that the self-trapped solution is an artifact of the DMF approach and does not exist also for large impurity-boson interactions.
Mean-field approaches to the Bose polaron in a frame of relative coordinates between bosons and impurity, obtained by a Lie-Low-Pines transformation, amended by a linearized fluctuation analysis, have been shown to provide accurate descriptions of Bose polarons for heavy impurities ($M/m\geq 3$) and weak boson-boson interactions $(\gamma\leq 1)$ [53, 11, 54, 12, 13]. We here showed that the same is true for strongly interacting bosons. To this end we calculated the polaron energy and derived analytical approximations for large but finite Tonks parameters $1\ll\gamma<\infty$ and arbitrary boson-impurity couplings by using the mapping to weakly interacting fermions.
Finally we numerically calculated the short-distance interaction potential between two impurities in Born-Oppenheimer approximation for arbitrarily strong boson-boson interactions. For small Tonks parameter $\gamma\leq 1$ we verified the results of [13] were a linear short-distance behavior was predicted. In the strong interaction limit $\gamma\gg 1$ we found oscillatory modulations in the potential in agreement with low-energy approximations [28, 29] and extending them to short distances, relevant for bi-polaron bound states.
Note: After submission of our manuscript a work by Gomez-Lozada et al. appeared studying polarons in 1D lattices of interacting bosons [55], also showing (among other things) that correlations prevent phase separation and self-trapping.
Acknowledgement
The authors gratefully acknowledge financial support by the DFG through SFB/TR 185, Project No.277625399. E.V.M. furthermore acknowledges support by Fundació Agustí Pedro Pons.
Appendix
VI.1 DMRG simulation of quantum impurity in a 1D Bose gas
In order to apply DMRG for the continuum model let us first look at the Hamiltonian where both bosons and the impurity are described in second quantization: